# AnomalyControl: Learning Cross-modal Semantic Features for Controllable Anomaly Synthesis

### v1

@@ 3.2.3 Semantic Guided Adapter
This setup allows the SGA to integrate anomaly-specific semantic signals effectively, ensuring high-quality and controllable anomaly synthesis.
Algorithm 1 AnomalyControl Training Algorithm
Data: visual anomaly descriptor $I_{a}$, textual anomaly descriptor $T_{a}$, anomaly region mask $M_{A}$, targeted text prompt $T_{t}$, time steps $T$
Input: Number of training steps $T_{train}$, Guidance steps $T_{g}$
Output: Trained model parameters for AnomalyControl
for $t_{train}\leftarrow 1$ to $T_{train}$ do
       Sample a batch of images $I$ from the training dataset $I_{a}$ Encode images to latent space: $z=\text{Encoder}(I)$ Sample noise $\epsilon\sim\mathcal{N}(0,1)$ Sample time step $t\sim\text{Uniform}(1,T)$ Add noise to the latent variable $z$ to obtain $z_{t}$ Set trainable attention guidance variable $e_{g}=0$ Extract text embedding $e_{t}$ and visual embedding $e_{v}$ from the frozen VLM. for $g\leftarrow 1$ to $T_{g}$ do
             Generate attention map $A=\text{VLM}_{\text{crossattn}}(e_{t},e_{v}+e_{g})$ Compute average attention map $\bar{A}_{\text{anomaly}}$ by Eq. 2 Compute $E(\bar{A}_{\text{anomaly}},M_{A})$ by Eq. 3 Update $e_{g}$ by Eq. 4
       end for
      Generate cross-modal semantic feature $C^{\prime}=\text{VLM}(e_{t},e_{v}+e_{g})$ Compute loss $\mathcal{L}$ by Eq. 6 Backpropagate $\mathcal{L}$ and update SGA parameters
end for
return Trained AnomalyControl model parameters
Table 2: Generation quantitative results with IS and IC-LPIPS (abbreviated as IL) on MVTec AD dataset. Bold values indicate the best results, and underlined values indicate the second-best results.
DiffAug [39] IS 1.59 1.72 1.34 1.19 1.96 1.67 2.07 1.58 1.53 1.10 1.93 1.33 1.34 2.05 1.30 1.58
CDC [22] IS 1.52 1.97 1.37 1.25 1.97 1.97 1.80 1.55 1.56 1.13 2.10 1.63 1.61 2.05 1.30 1.65
SDGAN [21] IS 1.57 1.89 1.49 1.18 1.95 1.85 2.04 1.45 1.61 1.17 2.53 1.78 1.76 2.12 1.25 1.71
DefGAN [35] IS 1.39 1.70 1.59 1.24 2.01 1.87 2.12 1.47 1.61 1.19 2.35 1.85 1.47 2.19 1.25 1.69
DFMGAN [8] IS 1.62 1.96 1.59 1.23 1.97 1.93 2.06 1.49 1.63 1.12 2.39 1.82 1.64 2.12 1.29 1.72
AnoDiff [13] IS 1.58 2.13 1.59 1.16 2.04 2.13 1.94 1.96 1.61 1.28 2.54 1.68 1.57 2.33 1.39 1.80
AnoXFusion [12] IS 1.62 2.13 1.68 1.16 2.12 2.24 2.00 1.96 1.63 1.29 2.55 1.80 1.64 2.09 1.37 1.82
IL 0.19 0.42 0.20 0.22 0.45 0.33 0.42 0.34 0.27 0.29 0.52 0.22 0.36 0.38 0.26 0.33
Ours IS 1.63 2.14 1.69 1.18 2.26 2.12 2.08 1.98 1.64 1.25 2.53 1.80 1.66 2.20 1.40 1.84
4 Experiments
Figure 3: Qualitative Evaluation. It is observed that our method could exhibit large improvements in realism and generalization. (a) Comparison with anomaly synthesis methods. Each row showcases anomalies across various object categories, where binary-value masks indicate the position of the generated anomaly regions. (b) Comparison with image generation methods. From left to right, the columns display a normal image with a mask indicating the target anomaly region, a visual anomaly descriptor to provide visual guidance, and the generated results using different methods (i.e., reference-only mode ControlNet, IP-Adapter, and our method), respectively.
4.1 Experimental setup
Datasets. We conducted our experiments on the widely used MVTec AD dataset [3, 4], which comprises 5,354 images across 15 categories for industrial anomaly detection tasks, including 10 object categories and 5 texture categories. Each category is represented by a set of defect-free training images and a test set containing images with various defects as well as defect-free images. Pixel-level annotations are provided for all existing anomalies in this dataset. We utilized the captions from MVTec AD Caption [12] for training and testing, which were carefully designed to ensure cross-modal complementarity. For data partitioning, we use the lowest one-third of ID numbers from the anomaly dataset as the training set, with the remaining two-thirds reserved for testing. For each anomaly type, 1,000 pairs of anomaly images and corresponding masks are generated using the mask generation method from AnoDiff [13]. These pairs are employed for downstream anomaly detection and localization tasks.
@@ 4.1 Experimental setup
Metrics. Following previous methods [13, 12], to evaluate the quality and diversity of synthesized anomalies, we used the Inception Score (IS) [30], which measures overall quality and diversity in generated images, and the Intra-cluster Pairwise Learned Perceptual Image Patch Similarity (IC-LPIPS) [23] to evaluate perceptual similarity and the generalization of synthetic anomalies. For evaluating accuracy in downstream tasks such as anomaly detection and localization, we employed measurements through pixel-level and image-level Area Under the Receiver Operating Characteristic Curve (AUROC), Average Precision (AP), and F1-max scores. Consistent with DRAEM[34]’s procedure, we generated 1,000 images per anomaly category and trained a U-Net alongside normal samples.
Implementation Details. AnomalyControl is implemented with the HuggingFace Diffusers library [24] and built on the Stable Diffusion 1.5 (SD1.5) model. The VLM utilized in CSM is based on BLIP-2 [16], which provides robust semantic feature extraction for cross-modal inputs. We incorporate a new image cross-attention layer into each of SD1.5’s 16 cross-attention layers, setting the cross-modal guidance strength parameter $\gamma=1$ for balanced guidance. The number of guidance steps $T_{g}$ in ASEA is set to 3. We use the AdamW optimizer [18] with a fixed learning rate of 0.0001 and weight decay of 0.01. For classifier-free guidance, text or cross-modal semantic features are independently dropped with a probability of 0.05, and both are dropped simultaneously with a probability of 0.05. Training is conducted on a single NVIDIA 4090 GPU, requiring approximately three days for 30K iterations. During inference, a 30-step DDIM sampler is used, with a guidance scale of 7.5, to achieve a balance between fidelity and generalization in the generated anomalies.
Baselines. For anomaly synthesis comparisons, we selected several recent high-performing methods as baselines, including DiffAug [39], CDC [22], SDGAN [21], DefGAN [35], DFMGAN [8], AnoDiff [13], and AnoXFusion [12]. To evaluate performance in downstream anomaly detection and localization tasks, we compared our approach with existing outperforming techniques such as DRAEM [34], PRN [36], DFMGAN [8], AnoDiff [13], and AnoXFusion [12].
4.2 Comparison in Anomaly Synthesis
@@
Table 2 illustrates the quantitative results for the generation task on the MVTec AD dataset. Our method achieves the highest average IS of 1.84 and the highest average IL of 0.35, surpassing all other methods in both overall quality and generalization. These results demonstrate our model’s ability to generate high-quality, diverse anomalies that closely resemble real-world anomalies while maintaining robustness across different anomaly types.
Figure 3 provides a qualitative evaluation of our approach’s effectiveness in anomaly synthesis. (a) In comparison with anomaly synthesis methods, our approach outperforms existing methods by producing anomalies with enhanced realism, closely aligning with real-world defect patterns. (b) In the comparison with image generation methods, we selected reference images from different categories than the target images (for the first three rows) or from the same category with anomalies in different locations (for the fourth row). This setup demonstrates our method’s flexibility in transferring anomalies across various contexts and positions, thereby enhancing the model’s generalization capability. Our approach demonstrates superior control over anomaly placement and appearance, accurately capturing the characteristics of the reference anomalies. This improvement in visual fidelity highlights our method’s advantage in generating realistic, high-quality anomalies.
4.3 Comparison in Downstream Tasks
@@
As shown in Table 3, our model consistently outperformed other anomaly synthesis methods across most evaluation metrics, demonstrating superior efficacy in both anomaly detection and localization. By enabling fine-grained controls over the appearance of anomalies, our approach can ensure that the generated samples closely resemble real-world defects. This capability is crucial in providing diverse and realistic training data, allowing downstream models to generalize better to real-world anomalies and improve both detection sensitivity and localization accuracy.
Table 3: Results of anomaly detection on MVTec AD dataset. P denotes the results of anomaly localization (pixel-level) and I denote the results of anomaly detection (image-level).
@@
DRAEM [34] 92.2 54.1 53.1 94.6 97.0 94.4
PRN [36] 96.9 66.2 64.7 91.6 96.6 92.4
DFMGAN [8] 90.0 62.7 62.1 87.2 94.8 94.7
AnoDiff [13] 99.1 81.4 76.3 99.2 99.7 98.7
AnoXFusion [12] 99.3 86.1 80.6 99.2 99.8 98.7
@@
Figure 4: Effect of guidance steps $T_{g}$ in ASEA. As $T_{g}$ increases, the loss value is decreased, where ASEA progressively refines VLM’s attention on anomaly regions. The generated images exhibit improved anomaly realism, confirming the effectiveness of ASEA in focusing on specified anomaly regions and producing more visually realistic anomalies.
4.4 Ablation Study
Table 4: Ablation Study. All modules (CSM, ASEA, and SGA) could improve the effectiveness of the method. We also study the impact of different guidance steps $T_{g}$ in ASEA.
@@ 4.4 Ablation Study
From the experimental results in Table 4, it is observed that the stepwise introduction of each module significantly improves model performance. First, with only the SGA module, IS and IL increase to 1.66 and 0.23, respectively, while AUC-P and AUC-I reach 92.7 and 95.4. At this stage, the model only uses the visual anomaly descriptor, functioning as an adapter based solely on the image prompt. The addition of the SGA module shows that SGA can strengthen the focus on the anomaly regions in the image. When the CSM module is further introduced, IS rises from 1.66 to 1.69, IL from 0.23 to 0.25, and AUC-P and AUC-I reach 93.1 and 95.8. This improvement indicates that the CSM module incorporates the textual anomaly descriptor, integrating it with the visual anomaly descriptor for cross-modal semantic fusion, thereby improving the quality and coherence of anomaly generation. As described in the methods section, the CSM module captures cross-modal semantic features of the anomaly region through multimodal interactions, resulting in anomaly features that align better with the intended descriptions. Finally, for the ASEA module, different guidance steps $T_{g}$ demonstrate its impact on model performance. When $T_{g}=3$, all metrics reach optimal values, providing the best balance between computational efficiency and performance. Increasing $T_{g}$ to 4 or 5 shows marginal improvements, indicating that performance gains plateau while computational costs continue to rise. As shown in Figure 4, ASEA progressively refines VLM’s focus on specified anomaly areas, leading to generated images with higher anomaly realism. This indicates ASEA’s effectiveness in concentrating attention, resulting in visually realistic and detailed anomalies. By enhancing focus on anomaly-specific regions, ASEA effectively helps to generate high-quality, detailed anomaly images.
5 Conclusion

### v4

@@ Methods
where $z_{m}$ represents the masked image latent. This setup allows SGA to integrate anomaly-specific semantic signals effectively, ensuring high-quality and controllable synthesis.
Experiments
Experimental Setup
Datasets. We conduct experiments on three datasets, including MVTec AD (Bergmann et al. 2019, 2021), MPDD (Jezek et al. 2021), and ViSA (Zou et al. 2022), to evaluate the performance of our method. Following the protocol in AnoDiff (Hu et al. 2024b), we adopt the same data split: the first one-third of the anomaly dataset IDs is used for training the anomaly synthesis model, and the remaining two-thirds for testing. The anomaly synthesis model generates 1,000 pairs of anomaly images and corresponding masks for each anomaly category using normal data. The masks are generated using the mask generation method from AnoDiff (Hu et al. 2024b). These pairs of anomaly images and their corresponding masks, along with normal images from the training set, are then used to train downstream anomaly detection models, following the same approach as DRAEM (Zavrtanik, Kristan, and Skočaj 2021) and AnoDiff (Hu et al. 2024b).
@@
Metrics. Inception Score (IS) (Salimans et al. 2016) (Hu et al. 2024b) is to evaluate the quality and diversity of synthesized anomalies. Intra-cluster Pairwise Learned Perceptual Image Patch Similarity (IC-LPIPS, abbreviated as IL) (Ojha et al. 2021b) is used to evaluate perceptual similarity and the generalization of synthetic anomalies. For downstream tasks such as anomaly detection and localization, we employed measurements through pixel-level and image-level Area Under the Receiver Operating Characteristic Curve (AUROC), Average Precision (AP), and F1-max scores.
Implementation Details. AnomalyControl is implemented with the HuggingFace Diffusers library (Patrick von Platen, Suraj Patil, Anton Lozhkov, Pedro Cuenca, Nathan Lambert, Kashif Rasul, Mishig Davaadorj, and Thomas Wolf 2022) and built on the Stable Diffusion 1.5 (SD1.5) model. The VLM utilized in CSM is based on BLIP-2 (Li et al. 2023), which provides robust semantic feature extraction for cross-modal inputs. We incorporate a new cross-attention layer into each of SD1.5’s 16 cross-attention layers, setting the cross-modal guidance strength parameter $\gamma=1$ for balanced guidance. The number of guidance steps $T_{g}$ in ASEA is set to 3. Additional details regarding the implementation are provided in the Appendix.
Figure 3: Display of anomaly synthesis on 15 objects from the MVTec AD dataset. The first column shows the normal image, the second column represents the mask area where anomalies need to be generated, the third column contains the visual anomaly descriptor providing visual guidance, and the fourth column displays the results generated by our method. Unlike simply copying the anomaly shown in the visual anomaly descriptor, our model predicts what the anomaly would look like in the given region indicated by the mask, considering the material and texture of the original normal image. The generated anomalies are consistent with the context of the image and are realistic in their appearance. Zoom in for better clarity and details.
Figure 4: Qualitative Evaluation. Our method could exhibit large improvements in realism and generalization compared with both anomaly synthesis and image generation methods. For example, in the grid and cable examples, our method produces the clearest and most realistic anomalies, closely matching the texture and material of the original image.
Figure 5: Effect of guidance steps $T_{g}$ in ASEA. As $T_{g}$ increases, the loss value decreases, where ASEA progressively refines VLM’s attention on anomaly regions. The generated images exhibit improved anomaly realism, confirming the effectiveness of ASEA in focusing on specified anomaly regions and producing more visually realistic anomalies.
Experimental Results
@@
Comparisons in Anomaly Synthesis. Table 2 illustrates the quantitative results for the generation task on the MVTec AD dataset. Our method achieves the highest average IS of 1.84 and the highest average IL of 0.35, surpassing all other methods in both overall quality and generalization. These results demonstrate our model’s ability to generate high-quality, diverse anomalies that closely resemble real-world anomalies while maintaining robustness across different anomaly types.
As shown in Figure 3, our approach can generate realistic anomalies that are contextually consistent with the original images. For instance, in the hazelnut example, the generated anomaly of a cracked and broken shell is highly consistent with the texture and appearance of the original object, while maintaining a precise match with the defined mask area. Similarly, in the capsule and wood examples, our method accurately generates anomalies like incorrect imprints and scratches, showing its capability to faithfully replicate defects.
@@
Comparisons in Downstream Tasks. As shown in Table 3, our model consistently outperformed other anomaly synthesis methods across most evaluation metrics. By enabling fine-grained controls over the appearance of anomalies, our approach can ensure that the generated samples closely resemble real-world defects. This capability is crucial in providing diverse and realistic training data, allowing downstream models to generalize better to real-world anomalies and improve both detection sensitivity and localization accuracy.
Qualitative Comparisons. As shown in Figure 4, our approach outperforms existing anomaly synthesis methods by producing anomalies with enhanced realism, closely aligning with real-world defect patterns. (1) For anomaly synthesis methods, each row showcases anomalies across various object categories, where binary-value masks indicate the position of the generated anomaly regions. Our method excels in generating realistic anomalies in regions where other methods fail to capture fine details, resulting in unrealistic anomalies, artifacts (e.g., DFMGAN and AnoDiff, “cable”), or misaligned/fail to generate corresponding anomalies within the mask (e.g., DFMGAN and AnoDiff, “grid”). (2) For image generation methods, from left to right, the columns display a normal image with a target mask, a visual anomaly descriptor to provide visual guidance, and the generated results using different methods (i.e., reference-only mode ControlNet, IP-Adapter, and our method), respectively. Our method generates anomalies that are well-aligned with the visual cues in the descriptor, producing defects with a high degree of realism and relevance.
MVTec AD
@@
DRAEM 92.2 54.1 53.1 94.6 97.0 94.4
PRN 96.9 66.2 64.7 91.6 96.6 92.4
DFMGAN 90.0 62.7 62.1 87.2 94.8 94.7
AnoDiff 99.1 81.4 76.3 99.2 99.7 98.7
@@
MPDD
Method AUC-P AP-P F1-P AUC-I AP-I F1-I
DRAEM 85.2 36.7 38.3 87.6 73.4 67.8
PRN 90.7 40.8 41.5 86.3 77.6 76.3
DFMGAN 92.4 41.6 45.5 92.5 78.5 78.9
AnoDiff 95.0 47.6 49.5 95.3 88.5 88.0
Ours 96.5 50.6 51.2 98.4 90.5 89.1
VisA
Method AUC-P AP-P F1-P AUC-I AP-I F1-I
DRAEM 88.8 25.4 24.6 84.1 88.2 87.4
PRN 87.0 20.8 21.5 80.5 81.9 82.3
DFMGAN 95.2 39.9 40.2 87.1 85.2 85.7
AnoDiff 98.0 33.2 37.1 96.2 96.9 92.6
Ours 99.0 42.5 48.0 96.8 97.2 93.5
Table 3: Results of anomaly detection. “-P” denotes the results of anomaly localization (pixel-level) and “-I” denote the results of anomaly detection (image-level). Bold values indicate the best results, and underlined values indicate the second-best results.
@@
Table 4: Ablation Study. All modules could improve the effectiveness of the method. We also study the impact of different guidance steps $T_{g}$ in ASEA. Bold values indicate the best results, and underlined values indicate the second-best results.
Effectiveness of ASEA. As shown in Figure 5, ASEA progressively refines VLM’s focus on specified anomaly areas, leading to higher anomaly realism for visually realistic and detailed anomalies. By enhancing focus on anomaly-specific regions, ASEA effectively helps to generate high-quality, detailed anomaly images.
@@
Ablation Study. As shown in Table 4, it is observed that all modules could improve the effectiveness of the method. With only the SGA module, IS and IL increase to 1.66 and 0.23, respectively, while AUC-P and AUC-I reach 92.7 and 95.4. At this stage, the model only uses the visual anomaly descriptor, functioning as an adapter based solely on the image prompt. The addition of the SGA module shows that SGA can strengthen the focus on the anomaly regions in the image. With the CSM module, IS rises from 1.66 to 1.69, IL from 0.23 to 0.25, and AUC-P and AUC-I reach 93.1 and 95.8. This improvement indicates that the CSM module incorporates the textual anomaly descriptor, integrating it with the visual anomaly descriptor for cross-modal semantic fusion, thereby improving the quality and coherence of anomaly generation. As described in the methods section, the CSM module captures cross-modal semantic features of the anomaly region through multimodal interactions, resulting in anomaly features that align better with the intended descriptions. With the ASEA module, different guidance steps $T_{g}$ demonstrate its impact on model performance. When $T_{g}=3$, all metrics reach optimal values, providing the best balance between computational efficiency and performance. Increasing $T_{g}$ to 4 or 5 shows marginal improvements, indicating that performance gains plateau while computational costs continue to rise.
We provide additional experimental results, analyses, and qualitative examples in the Appendix.
Conclusion